# Quantum Computing for All: Online Courses Built Around Interactive Visual Quantum Circuit Simulator

### v1

@@ Abstract
Since quantum computing is becoming a practical field, it requires students to gain hands-on experience with both quantum software and hardware. However, the learning curve for quantum computing should be smoother and not require deep programming expertise from the start. Further, even if there are already some quantum computers available for online access, the computing time may be limited, especially for students who are not enrolled in specialized programs. These reasons have motivated the University of Jyväskylä to develop an interactive quantum circuit simulator as a part of the existing learning platform.
The rest of paper is structured as follows, aligning with the phases of the Design Science Research (SDR) methodology [9]. We start by introducing the problem identification and motivation, followed by presenting the learning environment in which the solution is integrated. Then, we describe the quantum simulator implementation, and a demonstration on how it is used to conduct three specific tasks within the educational process. Subsequently, we discus how the visual editor meets the teaching goals for the target student audience. We finish with concluding remarks.
@@
Quantum computing is an interdisciplinary topic that cannot be put under a single scientific field. For that reason, also students from various study tracks are eager to learn about quantum computing. However, their different background and prerequisite knowledge poses an obvious challenge. A physics student knows about quantum mechanics, but may not be fluent in programming. A computer science student knows programming, but has not the needed theoretical background of quantum physics. Further, a third case is e.g. a student in business school, who understands how technology relates to business, and would hence keen to know the basics of quantum computing. These student personas, their educational background and expertise, as well as their expectations are summarized in Table 1. When developing the new teaching material, the different personas has served as imaginary student profiles and helped us to take into account their diverse viewpoints.
@@
At the started the project, TIM already had all the essential features for lecture-oriented course as well as for a MOOC. Our purpose was to design and build a tool that is functional in both types of courses. The quantum circuit simulator is implemented as a TIM plugin, and it can be customized to create different types of exercises. The teacher has options to customize the different components to be visible or not according to the context it is used in.
@@
The result part of the simulator interface provides several widgets that allow the visualization of the quantum computation. The bar chart widget shows the probability distribution of the output states. For larger circuits, rows where output probability is zero can be hidden from the chart to make the results easier to understand. The measurements widget allows the user to perform shot measurements and to visualize the sampled output value in a table containing the measurement number, and the corresponding input and output.
@@
The visibility of the elements of circuit simulator interface is configurable, therefore the user can first be shown a strip down version and then additional features can be gradually introduced making the simulator more approachable for new users.
@@
When there are a lot of simulation exercises on one web page the load time increases making it quite slow especially on slower devices like tablets. This was mitigates by lazy loading the simulator components. The simulator is fully rendered and the simulation computation is done when user clicks on the exercise.
The exercises can be automatically checked for correctness by specifying model circuit to compare the user’s circuit to. The circuit equality is checked by iterating over all $2^{nQubits}$ input combinations and checking that both circuit result in same probability vector. The number of checked inputs can be reduced by giving regex patterns to match each input bit-strings against and only running the simulation against these matching bit-string inputs.
@@
The quantum circuit simulator is a flexible tool to construct various exercises. In the following we demonstrate three different kind of exercises that are in used in our online courses.
@@
The action of a single quantum gate is simple. To achieve certain task, one may have to use a circuit composed of several gates. Their collective performance may be hard to infer. The circuit simulator offers a fantastic workaround as a student and experiment with gate constellations. The simulator gives an instant visualization of the output in the form of measurement outcome distribution.
@@
Once a student has reached some familiarity of quantum gates has been achieved, she can proceed to more interesting exercises. With the simulator the teacher can create arbitrary gates and label them in any desired way. This makes it possible to form exercises where a student has e.g. identify an unknown gate, or their combination, by using it in concatenation with other gates (see Figure 7). This kind of exercises are almost like entertaining puzzles. As the gate can be any of the previously introduced gates, a student necessarily has to review previously learned material. Some quantum gates, like the Z-gate, do nothing in the computational basis unless they are coupled with some other, appropriately chosen, quantum gates. This quantum feature makes additional excitement to suitably planned exercises.
@@ DISCUSSION
The spectrum of experiences that can be used by quantum technology practitioners (excluding hardware), can be divided into the following areas, as depicted in Figure 8: paper, paper-like, notebook, and integrated development environment (IDE). The paper is used mainly by practitioners with physics and mathematical background and does not necessitate computer systems to experimenting besides paper or a whiteboard. The paper-like provides a low threshold for using a computer system that offers an experience close to the paper, but with instant feedback aided by tools like TIM or IBM Quantum Composer. The notebook experience is familiar to scientist that are using Python programming language and quantum development toolkits (e.g., Qiskit, Cirq [^4] https://quantumai.google/cirq or Pennylane [^5] https://pennylane.ai), and use tools like Jupyter and IBM Quantum Lab to develop algorithms and perform experiments. The IDE experience is tailored for software developers that use advanced tools like Visual Studio Code [^6] https://code.visualstudio.com or PyCharm [^7] https://www.jetbrains.com/pycharm/. Among our target student groups, the paper-like experience provided by the TIM’s quantum circuit simulator offers the optimal experience, as it does not require familiarity neither with programming languages, nor mathematical concepts, instead leading students to experiment with circuits by drag-and-dropping gates and instantly see the results.
Since February 2024, we have been running the first MOOC for early adopters. These are mainly students from the IT faculty that are keen in learning quantum computing. Hence, from the student personas described earlier, we have not covered all, and it is too early to draw definite conclusions. We can, however, make some preliminary observations.
@@
Given the diverse backgrounds of the students for whom the MOOC is designed, we anticipated that the difficulty of tasks would vary among students. However, the additional hints provided after most tasks have effectively served their purpose, preventing any task from becoming a significant bottleneck. We estimate the difficulty of tasks by the average number of attempts students make before arriving at the correct answer. While this average exceeds five for some tasks, it has not led to a decrease in motivation. In fact, students appear to appreciate a certain level of challenge. Since the number of attempts does not impact the course grade, students are motivated by the opportunity to tackle challenges without the pressure of exam failure or low grades. In this regard, the MOOC format is functioning as intended, fostering a supportive environment for learning.
@@
We recall that the simulator’s primary advantage over similar products is its seamless integration with the existing TIM learning platform. It’s not confined to a single course but can be utilized across various courses, allowing for the easy generation of exercises with diverse difficulty levels. The integration with TIM provides access to all necessary tools for a university course, including discussion forums, straightforward registration, and grading. The value of these features has been demonstrated, and the simulator has fulfilled its intended purpose. The predominant question from students has been, “When will the next course be available?” This underscores the clear demand for flexible, self-paced courses at the university level.

### v2

@@ Abstract
Since quantum computing is becoming a practical field, it requires students to gain hands-on experience with both quantum software and hardware. However, the learning curve for quantum computing should be smoother and not require deep programming expertise from the start. Further, even if there are already some quantum computers available for online access, the computing time may be limited, especially for students who are not enrolled in specialized programs. These reasons have motivated the University of Jyväskylä to develop an interactive quantum circuit simulator as part of the existing learning platform.
The rest of paper is structured as follows, aligning with the phases of the Design Science Research (SDR) methodology [9]. We start by introducing the problem identification and motivation, followed by the learning environment in which the solution is integrated. Then, we present the quantum simulator implementation, and describe how it is used to conduct three specific tasks within the educational process. Next, we discuss how the visual editor meets the teaching goals for the target student audience. We finish by exploring the feedback received from the first group of users that completed the course in which the system was used.
@@
Quantum computing is an interdisciplinary topic that cannot be put under a single scientific field. For that reason, also students from various study tracks are eager to learn about quantum computing. However, their different background and prerequisite knowledge poses an obvious challenge. A physics student knows about quantum mechanics, but may not be fluent in programming. A computer science student knows programming, but has not the needed theoretical background of quantum physics. Further, a third case is e.g. a student in business school, who understands how technology relates to business, and would hence keen to know the basics of quantum computing. These student personas, their educational background and expertise, as well as their expectations are summarized in Table 1. When developing the new teaching material, the different personas have served as imaginary student profiles and helped us to take into account their diverse viewpoints.
@@
At the start of the project, TIM already had all the essential features for lecture-oriented course as well as for a MOOC. Our purpose was to design and build a tool that is functional in both types of courses. The quantum circuit simulator is implemented as a TIM plugin, and it can be customized to create different types of exercises. The teacher has options to customize the different components to be visible or not according to the context it is used in.
@@
The results part of the simulator interface provides several widgets that allow the visualization of the quantum computation. The bar chart widget shows the probability distribution of the output states. For larger circuits, rows where output probability is zero can be hidden from the chart to make the results easier to understand. The measurements widget allows the user to perform shot measurements and to visualize the sampled output value in a table containing the measurement number, and the corresponding input and output.
@@
The visibility of the elements of the circuit simulator interface is configurable, therefore the user can first be shown the basic version and then additional features can be gradually introduced making the simulator more approachable for new users.
@@
When there are a lot of simulation exercises on one web page the load time increases making it quite slow especially on slower devices like tablets. This was mitigated by lazy loading the simulator components. The simulator is fully rendered and the simulation computation is done when user clicks on the exercise.
The exercises can be automatically checked for correctness by specifying a model circuit to compare the user’s circuit to. The circuit equality is checked by iterating over all $2^{nQubits}$ input combinations and checking that both circuit result in same probability vector. The number of checked inputs can be reduced by giving the regular expressions patterns to match each input bit-strings against, and only running the simulation against these matching bit-string inputs.
@@
The quantum circuit simulator is a flexible tool to construct various exercises. In the following we demonstrate three different kind of exercises that are used in our online courses.
@@
The action of a single quantum gate is simple. To achieve certain task, one may have to use a circuit composed of several gates. Their collective performance may be hard to infer. The circuit simulator offers a fantastic workaround as the student is encouraged to experiment with gate constellations. The simulator gives an instant visualization of the output in the form of measurement outcome distribution.
@@
Once a student has reached some familiarity of quantum gates, she can proceed to more interesting exercises. With the simulator the teacher can create arbitrary gates and label them in any desired way. This makes it possible to form exercises where a student has to identify an unknown gate, or their combination, by using it in concatenation with other gates (see Figure 7). This kind of exercises are almost like entertaining puzzles. As the gate can be any of the previously introduced gates, a student necessarily has to review previously learned material. Some quantum gates, like the Z-gate, do nothing in the computational basis unless they are coupled with some other, appropriately chosen, quantum gates. This quantum feature makes additional excitement to suitably planned exercises.
@@ DISCUSSION
The spectrum of experiences that can be used by quantum technology practitioners (excluding hardware), can be divided into the following areas, as depicted in Figure 8: paper, paper-like, notebook, and integrated development environment (IDE). The paper is used mainly by practitioners with physics and mathematical background and does not necessitate computer systems for experimenting besides paper or a whiteboard. The paper-like provides a low threshold for using a computer system that offers an experience close to the paper, but with instant feedback aided by tools like TIM or IBM Quantum Composer. The notebook experience is familiar to scientist that are using Python programming language and quantum development toolkits (e.g., Qiskit, Cirq [^4] https://quantumai.google/cirq or Pennylane [^5] https://pennylane.ai), and use tools like Jupyter and IBM Quantum Lab to develop algorithms and perform experiments. The IDE experience is tailored for software developers that use advanced software development tools like Visual Studio Code [^6] https://code.visualstudio.com or PyCharm [^7] https://www.jetbrains.com/pycharm/.
As the visual appearance of the TIM circuit simulator was inspired by the IBM Quantum Composer, we can say that their core functionality is similar, allowing the users to design and execute quantum circuits using metaphors specific to graphic user interface, and are able to explore their essential characteristics with specialised widgets. However, as the purpose of each tool varies – TIM is a tool optimised to facilitate learning in a formal education environment, while Quantum Composer is a commercial tool designed to smooth the onboard process of developers into IBM’s quantum platform – we expect the latter to add new visual programming capabilities and integrations with advanced simulators and actual quantum hardware at a much higher pace. Nevertheless, by being familiarized with the same metaphors and core capabilities, TIM’s users will be able to easy transition to more advanced commercial tools.
User feedback
Figure 9: Participants breakdown by department affiliation
Since February 2024, we have been running the first MOOC for a group of 60 early adopters (see the details in Figure 9). The Quantum Computing Essentials course is the first in a series of three courses planned at University of Jyväskylä. The language used is Finnish, and participants receive 2 study credits upon completion. The largest group is formed of students from the Information Technology, and Mathematics and Natural Sciences faculties that study quantum computing as part of their curriculum. Another large segment was formed by the university staff, followed by students from faculties that do not mandate the study of quantum computing, and participants from the open university. The study group covers two from the three student personas described earlier (e.g. physicist and software engineer), allowing us to draw some preliminary observations.
@@
Given the diverse backgrounds of the students for whom the MOOC is designed, we anticipated that the difficulty of tasks would vary among students. In order to prevent any task from becoming a significant bottleneck, we included additional hints in the course material that effectively served their purpose. To measure the difficulty of a task, we have used the average number of attempts made by the student before reaching the correct answer. Since the number of attempts does not impact the course grade, students are motivated by the opportunity to tackle the task challenges without the pressure of exam failure or low grades. In this regard, the MOOC format is functioning as intended, fostering a supportive environment for learning.
@@
We recall that the simulator’s primary advantage over similar products is its seamless integration with the existing TIM learning platform. It’s not confined to a single course but can be utilized across various courses, allowing for the easy generation of exercises with diverse difficulty levels. The integration with TIM provides access to all necessary tools for a university course, including discussion forums, straightforward registration, and grading. The value of these features has been demonstrated, and the simulator has fulfilled its intended purpose. Among our target student groups, the paper-like experience provided by the TIM’s quantum circuit simulator offers the optimal experience, as it does not require familiarity neither with programming languages, nor mathematical concepts, instead leading students to experiment with circuits by drag-and-dropping gates and instantly see the results.